find the equation of the line passes through the given point and has the given slope. Express your answer in slope incept form. (8,0), m = -3
Answer by mananth(16946)

slope = -3
point(8,0)(x1,y1)
m= slope
Te equation of point slope form
y-y1=m(x-x1)
y-0=-3(x-8)
y=-3x+24